# Learning Fused State Representations for Control from Multi-View Observations

### v1

@@ Abstract
Multi-View Reinforcement Learning (MVRL) seeks to provide agents with multi-view observations, enabling them to perceive environment with greater effectiveness and precision. Recent advancements in MVRL focus on extracting latent representations from multiview observations and leveraging them in control tasks. However, it is not straightforward to learn compact and task-relevant representations, particularly in the presence of redundancy, distracting information, or missing views. In this paper, we propose Multi-view Fusion State for Control (MFSC), firstly incorporating bisimulation metric learning into MVRL to learn task-relevant representations. Furthermore, we propose a multiview-based mask and latent reconstruction auxiliary task that exploits shared information across views and improves MFSC’s robustness in missing views by introducing a mask token. Extensive experimental results demonstrate that our method outperforms existing approaches in MVRL tasks. Even in more realistic scenarios with interference or missing views, MFSC consistently maintains high performance.
reinforcement learning, multi-view learning
@@ 1 Introduction
In this work, we integrate the self-attention mechanism with bisimulation metric, offering a powerful approach to multi-view representation learning and fusion. Inspired by BERT (Devlin, 2018) and ViT (Dosovitskiy, 2020), a learnable fusion [state] embedding for MVRL tasks is introduced to avoid biases from fixed views, and dynamically capture the globally optimal fusion of multi-view information. By aligning the fusion of multi-view representations based on the similarity of task-relevant information, we enable robust multi-view integration and efficient control. However, bisimulation metric learning heavily depends on rewards to fuse state representations, which may cause it to overlook fine-grained information or interactions unique to each view that are not directly reflected in the reward signal. Secondly, bisimulation assumes complete and high-quality observations, making it vulnerable to scenarios where observations are incomplete, as it lacks mechanisms to compensate for missing views.
@@
[…]
@@
  • We propose a novel multi-view representation learning and fusion algorithm MFSC, which is the first to incorporate bisimulation into the MVRL framework. This integration enables the extraction of task-relevant representations from redundant multi-view observations. Our method can be combined with any existing RL algorithms with minimal adjustments.
@@ 2 Related Work
Multi-View Reinforcement Learning. Effective state representation learning in MVRL aims to transform high-dimensional observations into a compact latent space. Li et al. (2019) propose a VAE-based algorithm that minimizes Euclidean distance between states encoded from different views, assuming a consistent primary view. Keypoint3D (Chen et al., 2021) uses 3D reconstruction to learn keypoints from third-person views, requiring camera calibration. Lookcloser (Jangir et al., 2022) introduces a cross-view attention mechanism for aggregating egocentric and third-person representations without calibration, though at higher computational cost. F2C (Hwang et al., 2023) uses conditional variational information bottlenecks (CVIBs) to model state space, showing robustness to missing views. MV-MWM (Seo et al., 2023b), based on Dreamer-v2 (Hafner et al., 2021), applies dual masking and pixel-level reconstruction tasks during pretraining. Since MV-MWM relies on expert demonstrations, we did not include a direct comparison. MVD (Dunion & Albrecht, 2024) employs contrastive learning to disentangle multi-view representations for control tasks. For a detailed comparison of MFSC with other algorithms, refer to Table 1 and Appendix B.2.
@@ 3.1 Multi-View Markov Decision Processes
A Multi-View Markov Decision Process is defined as the following tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{\vec{O}},\mathcal{P},\Omega,\mathcal{R},\gamma\rangle$. $\mathcal{S}$ represents the set of states $s$ in the environment, $\mathcal{A}$ is a set of actions $a$, $\mathcal{\vec{O}}=\{\mathcal{O}^{k}\}_{k=1}^{K}$ represents the set of $K$ observations $\vec{o}=\{o^{k}\}_{k=1}^{K}$. With the assumption that each multi-view observation $\vec{o}\in\mathcal{\vec{O}}$ uniquely determines its generating state $s\in S$, we can obtain the latent state regarding its multi-view observation by a projection function $\phi\left(\vec{o}\right):\vec{\mathcal{O}}\rightarrow\mathcal{S}$. Therefore, $s$ and $\phi\left(\vec{o}\right)$ can be used interchangeably. $\mathcal{P}(s^{\prime}|s,a)=\Pr(s_{t+1}=s^{\prime}|s_{t}=s,a_{t}=a)$ is the transition dynamics distribution. The corresponding transition function under the multi-view observation space is defined $\vec{o}^{\prime}\sim\mathcal{\hat{P}}(\vec{o}^{\prime}|\vec{o},a)$, where $\hat{\mathcal{P}}\left(\vec{o}^{\prime}|\vec{o},a\right)=\Omega\left(\vec{o}^{\prime}|s^{\prime}\right)\mathcal{P}\left(s^{\prime}|s,a\right)$ and $\Omega\left(\vec{o}|s\right)=\prod_{k=1}^{K}{\text{Pr}}\left(o_{t}^{k}=o^{k}|s_{t}=s\right)$ is the probability distribution of joint observation. $\mathcal{R}(s,a)\in\mathbb{R}$ is the immediate reward function for taking action $a$ at state $s$, $\gamma\in\left[0,1\right)$ is the discount factor. The goal of the agent is to find the optimal policy $\pi(a|s)$ to maximize the expected reward: $E_{s_{0},a_{0},\dots}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$.
@@ 3.2 Bisimulation
Bisimulation is used to formalize the concept of state similarity. The core principle is that two states are deemed equivalent if their future behaviors, i.e. state transitions and rewards given the same actions, are indistinguishable under the same policy. Castro et al. (2021) introduced the MICo update operator as a mathematical tool for constructing a $\pi$-bisimulation metric. This metric evaluates state similarity by focusing exclusively on the actions dictated by a specific policy $\pi$, rather than considering the entire action space.
@@
Castro et al. (2021) Let $\mathcal{M}$ be the set of all measures over $S$. The MICo update operator $\mathcal{F}^{\pi}:\mathcal{M}\to\mathcal{M}$ is defined as:
@@
Castro et al. (2021) The MICo update operator $\mathcal{F}^{\pi}$ has a unique fixed point $g^{\pi}_{\sim}$.
@@ 4.1 Bisimulation Metric in Multi-View Setting
The aggregator $\phi$ maps multi-view observations into a more compact space of summarizations $\mathcal{Z}$, defined as: $\phi:\mathcal{O}^{1}\times\mathcal{O}^{2}\times\cdots\times\mathcal{O}^{k}\rightarrow\mathcal{Z}\in\mathbb{R}^{d}$. By clustering observations to be similar under the bisimulation metric, the original MVRL problem can be approximated as solving a latent MDP ¯$\bar{\mathcal{M}}=\langle\mathcal{Z},\mathcal{A},\bar{\mathcal{P}},\Omega,\bar{\mathcal{R}},\gamma,p_{0}\rangle$.
@@ 4.2 Latent MDP Construction Assumptions
There exists a latent space $\mathcal{Z}$ with dimension lower than that of $\mathcal{O}^{1}\times\mathcal{O}^{2}\times\cdots\times\mathcal{O}^{k}$, and a bounded approximation error $\eta>0$, such that the induced latent MDP ¯$\bar{\mathcal{M}}=\langle\mathcal{Z},\mathcal{A},\bar{\mathcal{P}},\Omega,\bar{\mathcal{R}},\gamma,p_{0}\rangle$ satisfies:
@@ 5.1 Self-Attention Fusion Module
In this section, we will introduce how bisimulation can be incorporated into multi-view representation learning and aggregation. Specifically, our method comprises two core submodules: (i) Convolutional Feature Embedding, generating embeddings of the original high-dimensional multi-view observations; (ii) Self-Attention Fusion Module, integrating multi-view representations based on bisimulation metric.
@@
The output after $L$ attention layers is $,z_{L}=\{z^{0}_{L},z^{1}_{L},$$\dots,z^{k}_{L}\}$, where $z^{0}_{L}$ represents the final state fusion embedding. To extract task-relevant representations from multi-view observations, bisimulation metric learning is integrated into the state fusion process. Specifically, the bisimulation metric for a given policy $\pi$ (Eq.2) employs the measurement $G$, as defined in SimSR (Zang et al., 2022). Here, $G$ is based on cosine distance, which offers lower computational complexity compared to the Wasserstein distance while effectively mitigating representation collapse.
@@
$\left\{\begin{array}[]{l}G_{\odot}=\left|r_{\vec{o}_{i}}^{\pi}-r_{\vec{o}_{j}}^{\pi}\right|+\gamma G_{\phi_{\omega}}(s_{i}^{\prime},s_{j}^{\prime})\\ s_{i}^{\prime}\sim\mathcal{P}\left(\cdot\left|\phi_{\omega}\right.\left(\vec{o}_{i}^{\prime}\right),a\right),s_{j}^{\prime}\sim\mathcal{P}\left(\cdot\left|\phi_{\omega}\left(\vec{o}_{j}^{\prime}\right),a\right.\right)\end{array}\right.$ (10)
@@ 6 Experiment
Through our experiments, we aim to investigate the following questions: (1) How does the performance of MFSC compare to existing MVRL algorithms? (2) Has MFSC successfully captured task-relevant representations? And, how does the Self-Attention Fusion Module operate across different views? (3) How does the MFSC perform in the presence of noise interference? (4) To what extent can MFSC handle tasks with missing views? (5) What is the relative importance and contribution of each part of MFSC?
@@ 6.2 Evaluation on Meta-World and PyBullet’s Ant
Figure 3: Performance on PyBullet’s Ant.
PyBullet’s Ant. As shown in Fig.3, the performance of our method in the ‘Ant’ environment is comparable to Keypoint3D. During training, our algorithm exhibits stable and consistent performance improvement. Even in the more practical ‘Ant No Color’ task, our method exhibits the highest performance, demonstrating its ability to effectively capture and aggregate task-relevant information from multi-view observations. In contrast, Keypoint3D exhibits a significant performance decline after losing key color annotations.
@@ 6.3 Visualization and Analysis
The visualization above demonstrates that MFSC can effectively extract and fuse task-relevant representations from multiple views. The fused representations enhance the performance of downstream RL tasks by offering a more robust and comprehensive understanding of the environment.
@@ 6.4 Robustness Against Noisy Views
Figure 6: Partially visualization of the learned representation on the noisy view.
@@
6.5 Robustness Against Missing Views
To explore whether MFSC has the ability to counteract missing views, we systematically evaluate the performance of MFSC under conditions of missing views across three tasks from Meta-World. During training, we explicitly introduced a mask token and leveraged the masking and latent reconstruction to learn shared representations across views. In the testing phase, we simulate real-world scenarios with incomplete view information by randomly omitting observations from one view and replacing the missing view with the mask token. Fig.LABEL:fig:missing_view summarizes the performance comparison between MFSC and two other competitive MVRL algorithms, LookCloser and F2C, under both missing view and full view conditions. The results indicate that, although missing views can affect certain task-specific details, such as object position or robotic arm posture, MFSC still demonstrates improved convergence speed and results across the majority of tasks. Compared to LookCloser, the performance degradation of MFSC is relatively limited across all tested tasks, and it even achieves performance comparable to the previous SOTA algorithm for missing views, F2C. This robustness to missing views is largely attributed to the introduction of multiview-based masking and latent reconstruction, which leverages cross-view dependencies to promote robust shared representation learning.
@@ 6.6 Ablation Study
Fig.LABEL:fig:ablation illustrates the returns of the algorithm across two benchmarks, Meta-World and PyBullet’s Ant, comparing three variations: MFSC (the proposed method), MFSC without bisimulation constraints (‘MFSC w/o Bis’), MFSC without Masking and Latent Reconstruction (‘MFSC w/o MLR’), and MFSC without Self-Attention Fusion Module (‘MFSC w/o SAFM’), which means an average of the embeddings for each view.
@@ 7 Discussion
Limitations and Future Work. One limitation of our approach lies in its reliance on masking techniques to handle missing views. For certain views containing crucial information, accurately reconstructing the true environmental state remains challenging, even when inferring from other observations. This is due to the incomplete complementarity of information between multiple views, particularly in scenarios involving complex state transitions or occlusions. To address this limitation, future research could explore the integration of state-space models (Lee et al., 2020) to better capture temporal dependencies, enabling more accurate state estimation even in the absence of certain views.

### v4

@@ Abstract
Multi-View Reinforcement Learning (MVRL) seeks to provide agents with multi-view observations, enabling them to perceive environment with greater effectiveness and precision. Recent advancements in MVRL focus on extracting latent representations from multiview observations and leveraging them in control tasks. However, it is not straightforward to learn compact and task-relevant representations, particularly in the presence of redundancy, distracting information, or missing views. In this paper, we propose Multi-view Fusion State for Control (MFSC), firstly incorporating bisimulation metric learning into MVRL to learn task-relevant representations. Furthermore, we propose a multiview-based mask and latent reconstruction auxiliary task that exploits shared information across views and improves MFSC’s robustness in missing views by introducing a mask token. Extensive experimental results demonstrate that our method outperforms existing approaches in MVRL tasks. Even in more realistic scenarios with interference or missing views, MFSC consistently maintains high performance. The project code is available at https://github.com/zpwdev/MFSC.
Machine Learning, ICML
@@ 1 Introduction
In this work, we integrate the self-attention mechanism with bisimulation metric, offering a powerful approach to multi-view representation learning and fusion. Inspired by BERT (Devlin, 2018) and ViT (Dosovitskiy, 2020), a learnable fusion [state] embedding for MVRL tasks is introduced to avoid biases from fixed views, and dynamically capture the optimal fusion of multi-view information. By aligning the fusion of multi-view representations based on the similarity of task-relevant information, we enable robust multi-view integration and efficient control. However, bisimulation metric learning heavily depends on rewards to fuse state representations, which may cause it to overlook fine-grained information or interactions unique to each view not directly reflected in the reward signal. Secondly, bisimulation assumes complete and high-quality observations, making it vulnerable to scenarios where observations are incomplete, as it lacks mechanisms to compensate for missing views.
@@
Keypoint3D (Chen et al., 2021)
Pixel Encoder-Decoder 3D keypoints
LookCloser (Jangir et al., 2022)
Encoder only Cross-attention
F2C (Hwang et al., 2023)
Encoder only CVIBs
MV-MWM (Seo et al., 2023b) Feature+View
Pixel Encoder-Decoder Self-attention
MVD (Dunion & Albrecht, 2024)
Encoder only Contrastive learning
MFSC (ours) Pixel
Latent Encoder only Self-attention+Bisimulation
@@
  • We propose a novel MVRL algorithm MFSC, which is the first to incorporate bisimulation into the MVRL framework. This integration enables the extraction of task-relevant representations from redundant multi-view observations. Our method can be combined with any existing RL algorithms with minimal adjustments.
@@ 2 Related Work
Multi-View Reinforcement Learning. Effective state representation learning in MVRL aims to transform high-dimensional observations into a compact latent space. Li et al. (2019) propose a VAE-based algorithm that minimizes Euclidean distance between states encoded from different views, assuming a consistent primary view. Keypoint3D (Chen et al., 2021) uses 3D reconstruction to learn keypoints from third-person views, requiring camera calibration. Lookcloser (Jangir et al., 2022) introduces a cross-view attention mechanism for aggregating egocentric and third-person representations without calibration, though at higher computational cost. F2C (Hwang et al., 2023) uses conditional variational information bottlenecks (CVIBs) to model state space, showing robustness to missing views. MV-MWM (Seo et al., 2023b), based on Dreamer-v2 (Hafner et al., 2021), applies dual masking and pixel-level reconstruction tasks during pretraining. MVD (Dunion & Albrecht, 2024) employs contrastive learning to disentangle multi-view representations for control tasks. MOSER (Wan et al., 2024) is a model-based approach that actively seeks the optimal perspective for learning task representations under multiple views to enhance performance. For a detailed comparison of MFSC with other closely related algorithms, refer to Table 1 and Appendix B.2.
@@ 3.1 Multi-View Markov Decision Processes
A Multi-View Markov Decision Process is defined as the following tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{\vec{O}},\mathcal{P},\Omega,\mathcal{R},\gamma\rangle$. $\mathcal{S}$ represents the set of states $s$ in the environment, $\mathcal{A}$ is a set of actions $a$, $\mathcal{\vec{O}}=\{\mathcal{O}^{k}\}_{k=1}^{K}$ represents the set of $K$ observations $\vec{o}=\{o^{k}\}_{k=1}^{K}$. With the assumption that each multi-view observation $\vec{o}\in\mathcal{\vec{O}}$ uniquely determines its generating state $s\in S$, we can obtain the latent state regarding its multi-view observation by a projection function $\phi\left(\vec{o}\right):\vec{\mathcal{O}}\rightarrow\mathcal{S}$. Therefore, $s$ and $\phi\left(\vec{o}\right)$ can be used interchangeably. $\mathcal{P}(s^{\prime}|s,a)=\Pr(s_{t+1}=s^{\prime}|s_{t}=s,a_{t}=a)$ is the transition dynamics distribution. The corresponding transition function under the multi-view observation space is defined $\vec{o}^{\prime}\sim\mathcal{\hat{P}}(\vec{o}^{\prime}|\vec{o},a)$, where $\hat{\mathcal{P}}\left(\vec{o}^{\prime}|\vec{o},a\right)=\Omega\left(\vec{o}^{\prime}|s^{\prime}\right)\mathcal{P}\left(s^{\prime}|s,a\right)$ and $\Omega\left(\vec{o}|s\right)=\prod_{k=1}^{K}{\text{Pr}}\left(o_{t}^{k}=o^{k}|s_{t}=s\right)$ is the probability distribution of joint observation. $\mathcal{R}(s,a)\in\mathbb{R}$ is the immediate reward function for taking action $a$ at state $s$, $\gamma\in\left[0,1\right)$ is the discount factor. The goal of the agent is to find the optimal policy $\pi(a|s)$ to maximize the expected reward: $\mathbb{E}_{s_{0},a_{0},\dots}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$.
@@ 3.2 Bisimulation
Bisimulation is used to formalize the concept of state similarity. The core principle is that two states are deemed equivalent if their future behaviors, i.e. state transitions and rewards given the same actions, are indistinguishable under the same policy. Conventional bisimulation metric needs to compute the Wasserstein distance over the transition distributions across all actions, which is computationally expensive. Instead, Castro (2020) developed $\pi\text{-bisimulation metric}$ which removes the requirement of considering all actions and only needs to consider the actions induced by a policy $\pi$. Castro et al. (2021) introduced the MICo update operator as a mathematical tool for constructing a $\pi\text{-bisimulation metric}$. This metric evaluates state similarity by focusing exclusively on the actions dictated by a specific policy $\pi$, rather than considering the entire action space.
@@
($\pi\text{-bisimulation metric}$ (Castro, 2020)). Let $\mathcal{M}$ be the set of all measures over $S$. The operator $\mathcal{F}^{\pi}:\mathcal{M}\to\mathcal{M}$ is defined as:
@@
Castro (2020) The operator $\mathcal{F}^{\pi}$ has a unique fixed point $g^{\pi}_{\sim}$ and $g^{\pi}_{\sim}$ is a $\pi\text{-bisimulation metric}$.
@@ 4.1 Bisimulation Metric in Multi-View Setting
The aggregator $\phi$ maps multi-view observations into a more compact space of summarizations $\mathcal{Z}$, defined as: $\phi:\mathcal{O}^{1}\times\mathcal{O}^{2}\times\cdots\times\mathcal{O}^{k}\rightarrow\mathcal{Z}\in\mathbb{R}^{d}$. By clustering observations to be similar under the bisimulation metric, the original MVRL problem can be approximated as solving a latent MDP $\bar{\mathcal{M}}=\langle\mathcal{Z},\mathcal{A},\bar{\mathcal{P}},\Omega,\bar{\mathcal{R}},\gamma,p_{0}\rangle$.
@@ 4.2 Latent MDP Construction Assumptions
There exists a latent space $\mathcal{Z}$ with dimension lower than that of $\mathcal{O}^{1}\times\mathcal{O}^{2}\times\cdots\times\mathcal{O}^{k}$, and a bounded approximation error $\eta>0$, such that the induced latent MDP $\bar{\mathcal{M}}=\langle\mathcal{Z},\mathcal{A},\bar{\mathcal{P}},\Omega,\bar{\mathcal{R}},\gamma,p_{0}\rangle$ satisfies:
@@ 5.1 Self-Attention Fusion Module
In this section, we will introduce how bisimulation can be incorporated into multi-view representation learning. Specifically, our method comprises two core submodules: (i) Convolutional Feature Embedding, generating embeddings of the original high-dimensional multi-view observations; (ii) Self-Attention Fusion Module, integrating multi-view representations based on bisimulation metric.
@@
The output after $L$ attention layers is $z_{L}=\{z^{0}_{L},z^{1}_{L},$ $\dots,z^{k}_{L}\}$, where $z^{0}_{L}$ represents the final state fusion embedding. To extract task-relevant representations from multi-view observations, bisimulation metric learning is integrated into the state fusion process. Specifically, the bisimulation metric for a given policy $\pi$ (Eq.2) employs the measurement $G$, as defined in SimSR (Zang et al., 2022). Here, $G$ is based on cosine distance, which offers lower computational complexity compared to the Wasserstein distance while effectively mitigating representation collapse.
@@
[…]
@@ 6 Experiment
Through our experiments, we aim to investigate the following questions: (1) How does the performance of MFSC compare to existing MVRL algorithms? (2) Has MFSC successfully captured task-relevant representations? And, how does the Self-Attention Fusion Module operate across different views? (3) To what extent can MFSC handle tasks with missing views? (4) How does the MFSC perform in the presence of noise interference? (5) What is the relative importance and contribution of each part of MFSC?
@@ 6.2 Evaluation on Meta-World and PyBullet’s Ant
PyBullet’s Ant. As shown in Fig.3, the performance of our method in the ‘Ant’ environment is comparable to Keypoint3D. During training, our algorithm exhibits stable and consistent performance improvement. Even in the more practical ‘Ant No Color’ task, our method exhibits the highest performance, demonstrating its ability to effectively capture and fuse task-relevant information from multi-view observations. In contrast, methods based on contrastive learning and pixel-level reconstruction tend to overemphasize local variations, making it difficult to capture task-relevant information and leading to significant performance degradation when key color cues are removed.
@@ 6.3 Visualization and Analysis
The visualization above demonstrates that MFSC can effectively extract and fuse task-relevant representations from multiple views. The fused representations enhance the performance of downstream RL tasks by offering a more robust and comprehensive understanding of the environment. By integrating information from diverse observational sources, MFSC enables more efficient policy learning and decision-making in complex control scenarios.
(a) (a) Performance comparison of MFSC under full-view and missing-view. The black and gray lines represent the final convergence results of F2C. (b) Performance of ablation study.
6.4 Robustness Against Missing Views
To explore whether MFSC has the ability to counteract missing views, we systematically evaluate the performance of MFSC under conditions of missing views across three tasks from Meta-World. During training, we explicitly introduced a mask token and leveraged the masking and latent reconstruction to learn shared representations across views. In the testing phase, we simulate real-world scenarios with incomplete view information by randomly omitting observations from one view and replacing the missing view with the mask token. Fig.LABEL:fig:missing_view summarizes the performance comparison between MFSC and two other competitive MVRL algorithms, LookCloser and F2C, under both missing view and full view conditions. The results indicate that, although missing views can affect certain task-specific details, MFSC still exhibits decent performance in some tasks. Compared to LookCloser, the performance degradation of MFSC is relatively limited, and it even achieves performance comparable to the previous SOTA algorithm for missing views, F2C. This robustness to missing views is largely attributed to the introduction of multiview-based masking and latent reconstruction, which leverages cross-view dependencies to promote robust shared representation learning.
@@ 6.5 Robustness Against Noisy Views
Figure 7: Partial visualization under the noisy view.
@@ 6.6 Ablation Study
Fig.LABEL:fig:ablation illustrates the returns of the algorithm across two benchmarks, Meta-World and PyBullet’s Ant, comparing three variations: MFSC, MFSC without bisimulation constraints (‘MFSC w/o Bis’), MFSC without Masking and Latent Reconstruction (‘MFSC w/o MLR’), and MFSC without Self-Attention Fusion Module (‘MFSC w/o SAFM’), which means an average of the embeddings for each view.
@@ 7 Discussion
Limitations and Future Work. One limitation of our approach lies in its reliance on masking techniques to handle missing views. For certain views containing crucial information, accurately reconstructing the true environmental state remains challenging. This is due to the incomplete complementarity of information between multiple views, particularly in scenarios involving complex state transitions. To address this limitation, future research could explore the integration of state-space models (Lee et al., 2020) to better capture temporal dependencies, enabling more accurate state estimation even in the absence of certain views.
@@
Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable feedback. This work was partially supported by the NSFC under Grants 92270125 and 62276024, and by the Fundamental Research Funds for the Central Universities, JLU under Grant 93K172025K01.